Opened 11 years ago

Closed 9 years ago

Last modified 7 weeks ago

#191 closed flight processing (fixed)

IPY07/09, flight day 249a/2008, Krafla

Reported by: jatt Owned by: benj
Priority: alpha 4 high Milestone: 2008 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by mark1)

Data location: ~arsf/arsf_data/in_progress/2008/flight_data/ipy/IPY07_09-2008_249a_krafla

Data arrived via SATA disk K on 16th September 2008.

Scientific details: see ~/arsf/arsf_data/in_progress/2007/ARSF_Applications/ipy07/IPY07_-_09_Wright_-_Application_and_Science_Case.pdf

PI: Dr Tim Wright

There is no base station data.

Sensors:

  • ATM
  • Eagle
  • Hawk
  • Lidar

Quicklook

TBD

Change History (17)

comment:1 Changed 11 years ago by mark1

  • Description modified (diff)

comment:2 Changed 11 years ago by jatt

  • Owner set to jatt

comment:3 Changed 10 years ago by mggr

  • Type changed from task to flight processing

comment:4 Changed 10 years ago by jatt

Processed lidar data downloaded to:-

~arsf/arsf-data/in_progress/2008/flight_data/ipy/IPY07_09-2008_249a_Krafla/lidar/ 08249a_Krafla/

comment:5 Changed 10 years ago by jatt

  • Owner changed from jatt to mark1

Delivery directory created at ~jatt/scratch/ARSF/IPY07_09-2008_249a_krafla

comment:6 Changed 10 years ago by mark1

The screenshot a249a133b.jpg shows abnormal characteristics.

Checking complete!

comment:7 Changed 10 years ago by mark1

  • Owner changed from mark1 to jatt

Delivery check was done by mare and mark1. Data has not been rsynced over (pending correction above).

comment:8 Changed 10 years ago by jatt

Replaced screenshot a249a133b.jpg, with one no longer showing abnormal characteristics.

Rsynced to /users/rsg/arsf/arsf_data/in_progress/2008/flight_data/ipy/IPY07_09-2008_249a_krafla

comment:9 Changed 10 years ago by jatt

All data sent 16/12/08

comment:10 Changed 10 years ago by benj

  • Component changed from Processing: general to Archiving
  • Owner changed from jatt to benj

Requested archiving by DTG on 28/09/09

comment:11 Changed 9 years ago by benj

DTG archiving finished 22/12/2009, data copied to Larsen and deleted off Bowhead to recover space. Still needs to be tarballed up and sent to NEODC.

comment:12 Changed 9 years ago by benj

  • Resolution set to fixed
  • Status changed from new to closed

Made available for NEODC to download 11/01/2010

comment:13 Changed 4 years ago by tec

Reprocessing Delivery Check
Project folder has an a in it but cant find ticket so its going here, also delivery and readme have a b as the flight letter, is this correct??

/users/rsg/arsf/arsf_data/2008/flight_data/ipy/IPY07_09-2008_249a_krafla/processing/delivery/IPY07_09-249b-hyperspectral-20140225/

Ran unix2dos
unix2dos: converting file ./sensor_FOV_vectors/hawk_viewvector_list.txt to DOS format...
unix2dos: converting file ./sensor_FOV_vectors/eagle_viewvector_list.txt to DOS format...

Missing files:
doc/ISO19115-stylesheet.xsl
doc/ISO19115-css.css

Level 1b
Eagle
Some dropped frames on the eagle lines, Should be mentioned in the readme.
6,7,8,9,10,11 have underflows on around the last 5-10 bands

Hawk
1,2,3,4,5,6,7,8,910,11,12,13 first 8 blank, last 9 blank, 75-90, 140-180,220-246 lots underflows

comment:14 Changed 4 years ago by dac

Delivery

Delivered APL processed Eagle and Hawk data to Janice Hunter & Alasdair Mac Arthur via FTP on 12/05/2015

comment:15 Changed 8 weeks ago by wja

Archiving
Old hyperspectral delivery (AZGCORR format) moved to previous_deliveries/.
JSONs created and added to PostGIS.
Queued to RSYNC to CEDA.

comment:16 Changed 8 weeks ago by wja

Archiving
Started RSYNC to CEDA.

comment:17 Changed 7 weeks ago by wja

Archiving
RSYNC to CEDA complete (4/3/19).

Note: See TracTickets for help on using tickets.